OpenStudy (anonymous):

in a public opinion poll, 624 people from a sample of 1100 indicated they would vote for a specific candidate. assuming this poll to be a correct indicator of the electorate, how many votes can the canidate expect to receive from a populatoin of 40,000?

OpenStudy (ciarán95):

If we are assuming that the pole is a correct indicator of the electorate/population, then we are assuming that the proportion of the electorate who said they would vote, out of the 1100 surveyed, for the candidate will be the same as the proportion of the 40000 who will actually vote for the candidate. So, if 624 out of the 1100 said they would vote for this specific candidate, then: 624/1100 = 56.72% of the polled electorate would vote for them. So, we assume that 56.72% of the whole 40000 electorate will vote for them come election time. 100% of the electorate/population= 40000 1% of the electorate = 40000/100 This means that 56.72% of the electorate = (40000/100) x 56.72, which you should now be able to work out to give you the number of votes the candidate can expect./